OpenStudy (anonymous):

An irregular object with a mass of 22 g is dropped into a graduated cylinder containing 50 mL of water. When the object is dropped in, the water level in the graduated cylinder rises to 53.5 mL. What is the density of the object?

OpenStudy (anonymous):

for solving any physics question: 1) you should underline the thing that is asked. Here: Density of the object 2) now search for the easiest related formula to the subject. Here: \[\rho=\frac{ m }{ V}\] which means:\[Density=\frac{ mass }{ Volume }\] 3) Check if needed data are available or not. Here needed data are: Mass and Volume and they are both mentioned in the question. Mass=22g Volume=53.5-50=3.5mL (The graduated cylinder showed 3.5 mL difference after you dropped that object in the water. So this difference is exactly because of that object. It is its Volume.) 4) BEFORE YOU SUBSTITUTE data in formula, check if the units are proper or not. Here the unit of mass is gram. The best-known unit of Volume that can be used beside grams is L. It's 3.5 milliliter and you should convert milliliter to Liter. It's a very easy conversion like this: \[3.5 mL=3.5\times 10^{-3}L\] 5)Now substitute data in formula and it'll be done: \[\rho=\frac{ m }{ v }=\frac{ 22g }{ 0.0035L }=6285\frac{ g }{ L }\] Do Great!

OpenStudy (anonymous):

If it's a PURE material it can be something like Lanthanum. Although Lanthanum reacts with water fast.